what is the best method to solve 4x+8y=20
-4x+2y=-30

Answer by Gogonati(855)

The best method to solve this system linear equations is adding both sides:
4x+8y=20
-4x+2y=-30
10y=-10
y=-1
Substitute y=-1 in the first equation 4x+8(-1)=20 ,solving this equation we find x=7 The ordered pair (7,-1) is the solution for our system.